OpenStudy (anonymous):

Determine the surface-area-to-volume ratio for a rectangular prism with the given dimension... 6. 5x1x1 7.10x1x1 8.5x5x5 9.10x10x10

OpenStudy (jmark):

Surface area of rectangular prism = 2(lh + lw + wh) Volume of rectangular prism = lwh 6. Let l = 5, w = 1 and h= 1 SA = 2(5 + 5+ 1) = 22 V = 5 SA/V = 22/5 7. Let l = 10, w = 1 and h= 1 SA = 2(10 + 10+ 1) = 42 V = 10 SA/V = 42/10 = 21/5 8. Let l = 5, w = 5 and h= 5 SA = 2(25 + 25+ 25) = 150 V = 125 SA/V = 150/125 = 6/5 9. Let l = 10, w = 10 and h= 10 SA = 2(100 + 100+ 100) = 600 V = 1000 SA/V = 600/1000 = 3/5 ICSE board papers - http://scoremore.beep.com/icse-board-papers-2014-07-18-1.htm

OpenStudy (anonymous):

thank you @JMark